a metal crystallizes into two cubic phases,fcc and bcc.whose unit cell lengths are 4.0 and 3.0 angstroms respectively.the ratio of densities of fcc and bcc is

Grade:

1 Answers

Anurag Kishore
14 years ago

Hi,

Use the eqn

$ = ZM/(a3 Na)

$ = density

z = rank (no. of atoms per unit cell)

a = unit cell length

Na = Avogadro number

$1/$2 = z1 a23 / z2 a13

         = 4 * 33 / 2 * 43

          = 27 / 32

 

because there are 4 atoms in fcc and 2 atoms in bcc

Thanks

Anurag